# Fine-Tuning Code Language Models to Detect Cross-Language Bugs

### v1

@@ 4.2.2. Performance of Small and Large CodeLMs
As shown in the results of Table 4, the impact of fine-tuning varies across different CodeLMs. We further explored the differences in fine-tuning effects based on the CodeLM’s size and complexity. According to the definition in (Zhang et al., 2024b), models with fewer than 1B parameters are categorized as small LMs, while models with 1B parameters or more are considered as large LMs.
@@
Small CodeLMs: Small-scale CodeLMs, such as UniXcoder-base, GraphCodeBERT-base, CodeBERT-base, CodeT5p-220M, and NatGen, exhibited significant performance improvements after fine-tuning. In particular, UniXcoder-base demonstrated the best overall performance, suggesting that small CodeLMs can gain more knowledge transfer and performance enhancement from fine-tuning on the CLB dataset. On the other hand, among the small CodeLMs, CodeT5p-770M, with 770M parameters, showed the poorest overall performance. Despite having the best recall metric among all models (0.8779), it performed the worst in the other four metrics. The result suggests that, for the CLB detection task, large CodeLMs do not necessarily outperform small CodeLMs in terms of overall performance.
Large CodeLMs: Among the large CodeLMs that we fine-tuned, apart from the three CodeLMs with 7B parameters, the other three large CodeLMs have 3B or fewer parameters, specifically, Deepseek-Coder-1.3B-base, Granite-3B-Code-base-2k, and StarCoder2-3B. These CodeLMs also showed notable improvements after fine-tuning, but their performance gains were somewhat smaller compared to the small CodeLMs. As for the remaining three 7B large models - CodeQWen1.5-7B, Qwen2.5_Coder-7B, and CodeLlama-7b-hf — the performance improvement after fine-tuning was relatively modest, with some metrics even showing a decline rather than an improvement.
Finding 1: Among all the 13 selected CodeLMs, UniXcoder-base and GraphCodeBERT-base demonstrated the most significant improvements after fine-tuning. Notably, UniXcoder-base is ranked first in all metrics, indicating that its understanding of code structure and semantics during the pre-training phase has reached a high level, allowing it to achieve substantial performance gains in the CLB detection task. Although most CodeLMs showed varying degrees of improvement after fine-tuning, the performance gains of Qwen2.5_Coder-7B and CodeLlama-7b-hf were relatively limited. In terms of model parameter size, small CodeLMs generally outperform large CodeLMs in overall performance.
4.3. RQ2: Performance of CodeLMs Fine-Tuned on Single-Language Bug Datasets for CLB Detection
Table 5. Details of the CodeNet
Language Category Count* Sum
Java Buggy 305,759 624,479
Clean 318,720
Python Buggy 1,387,527 2,959,324
Clean 1,571,797
* Only count the results after deduplication.
Table 6. Details of the CVEFixes (the part of Python and Java)
Language Category Count Sum
Java Buggy 3,102 8,387
Clean 5,285
Python Buggy 2,775 7,568
Clean 4,793
Tables 5 and 6 provide the specific details of Java and Python code in the CodeNet and CVEFixes datasets, respectively. The following are the methods we used to process these two datasets. Since our CLB dataset is balanced (with equal numbers of bug and clean code samples), we processed the CVEFixes and CodeNet datasets to ensure fair experimental results. For the CVEFixes dataset, we used all bug samples. As there are more clean samples than bug samples, we randomly selected an equal number of clean samples to match the bug count, resulting in 3,102 pairs of Java bug and clean samples, and 2,775 pairs of Python bug and clean samples. Given the large size of the CodeNet dataset, which contains over 3 million Java and Python code samples, utilizing the entire dataset for fine-tuning would be computationally expensive and impractical. To ensure consistency with our CLB dataset, which contains 11,126 samples, we constructed a subset of the CodeNet for our fine-tuning experiments. Specifically, we first balanced the dataset by randomly selecting an equal number of clean samples to match the number of bug samples. Subsequently, we combined Java and Python samples by class (i.e., merging Java bug samples with Python bug samples, and Java clean samples with Python clean samples), and then randomly selected 10,000 samples each from the bug and clean data to create the final dataset for fine-tuning.
@@ 4.3. RQ2: Performance of CodeLMs Fine-Tuned on Single-Language Bug Datasets for CLB Detection
Table 7. Performance comparison of CodeLMs on single-language bug vs. CLB detection after fine-tuning on the CodeNet dataset
CodeLM Evaluate in CodeNet Evaluate in our CLB Dataset
@@
CodeBERT-base 0.6605 0.6341 0.7590 0.6909 0.7214 0.5009 0.5006 0.7253 0.5924 0.4869
GraphCodeBERT-base 0.6620 0.6716 0.6340 0.6523 0.7206 0.5108 0.5081 0.6732 0.5792 0.5012
UniXcoder-base 0.6745 0.6480 0.7640 0.7012 0.7290 0.4973 0.4982 0.7343 0.5936 0.5028
CodeT5p-220M 0.6615 0.6525 0.6910 0.6712 0.7113 0.5162 0.5105 0.7864 0.6191 0.5072
NatGen 0.6430 0.6525 0.6120 0.6316 0.6867 0.4982 0.4987 0.6934 0.5802 0.5010
Table 8. Performance comparison of CodeLMs on single-language bug vs. CLB detection after fine-tuning on the CVEFixes dataset
CodeLM Evaluate in CVEFixes Evaluate in our CLB Dataset
@@
CodeBERT-base 0.6352 0.6000 0.8112 0.6898 0.6867 0.5036 0.5022 0.8366 0.6276 0.5024
GraphCodeBERT-base 0.6429 0.6277 0.7024 0.6629 0.6941 0.4937 0.4959 0.7684 0.6028 0.4999
UniXcoder-base 0.6480 0.6033 0.8639 0.7105 0.7021 0.5018 0.5011 0.7864 0.6122 0.4901
CodeT5p-220M 0.6165 0.6049 0.6718 0.6366 0.6815 0.5018 0.5012 0.7504 0.6010 0.5053
NatGen 0.6250 0.5904 0.8163 0.6852 0.6728 0.5152 0.5094 0.8217 0.6289 0.4938
Single-language Bug Dataset Validation Results after Fine-tuning: According to Tables 7 and 8, after fine-tuning, the five models show strong performance in detecting single-language bugs. For example, UniXcoder-base outperforms the others on both datasets, especially in recall (0.8639) and F1 score (0.7105) on the CVEFixes dataset, demonstrating its ability to efficiently identify most bug instances. Although the fine-tuning results of CodeT5p-220M on CVEFixes and Natgen on CodeNet are not as strong as those of the other models, they still show commendable performance. After fine-tuning, almost all models achieve F1 and AUC scores around 0.7, indicating excellent single-language bug detection performance. In terms of recall, seven models exceed 0.7, with CodeBERT-base and UniXcoder-base achieving recall rates above 0.8 on CVEFixes. While precision is not as strong as recall, it still demonstrates solid performance. These results indicate that CodeLMs fine-tuned on existing SOTA single-language bug detection datasets perform exceptionally well in a single-language environment, accurately identifying bugs.
Cross-language Bug Dataset Validation Results after Fine-tuning: However, despite the strong performance of the models on single-language bug datasets, their performance significantly drops when applied to the CLB dataset we collected. For example, almost all models show a drastic decrease in AUC to around 0.5 on both the CVEFixes and CodeNet datasets, indicating that their performance is essentially at the level of random guessing. Although some CodeLMs achieve recall rates above 0.7, even over 0.8, and F1 scores above 0.6, both accuracy and precision remain around 0.5.
Finding 2: Although CodeLMs fine-tuned on single-language bug datasets like CVEFixes and CodeNet perform well in single-language bug detection, they struggle with CLB detection, showing low F1 scores and AUC values. CLBs, caused by interactions between different PLs, differ fundamentally from single-language bugs. Results demonstrate that CodeLMs trained on single-language bug dataset cannot effectively detect CLBs, highlighting the necessity of specialized datasets and techniques to address the unique challenges of CLB detection.
@@ 4.4.2. Impact of the Token Sequence Length
Finding 3: Increasing the size of the fine-tuning dataset for CodeLMs can indeed improve the model’s performance in CLB detection to some extent. However, increasing the token sequence length does not necessarily lead to better performance. Different CodeLMs show varying sensitivity to token sequence length; some CodeLMs (e.g., CodeT5p-220M) perform exceptionally well with shorter token sequence lengths, and their performance decreases as the token sequence length increases. On the other hand, some CodeLMs (e.g., GraphCodeBERT-base) perform better with longer token sequence lengths.
@@ 4.5. RQ4: Impact of Code Comments on Fine-tuned CodeLM Performance
We have already conducted fine-tuning experiments and analysis on the CLB dataset without code comments, and the results are shown in Table 4. To answer RQ4, we carried out fine-tuning experiments on the CLB dataset with code comments, and the results are shown in Table 9. By comparing the performance of CodeLMs fine-tuned on the CLB dataset with and without comments, we analyze the impact of code comments on the performance of CodeLMs in the following two aspects.
Table 9. Performance of CodeLMs in detecting CLBs before and after fine-tuning on the CLB dataset with code comments
CodeLM Performance Before Fine-Tuning Performance After Fine-Tuning
Acc Prec Rec F1 AUC Acc Prec Rec F1 AUC
CodeBERT-base 0.4866 0.4903 0.6750 0.5680 0.4854 0.6304 0.6008 0.7768 0.6776 0.7108
GraphCodeBERT-base 0.5054 0.5195 0.1429 0.2241 0.5038 0.6604 0.6304 0.7754 0.6954 0.7898
UniXcoder-base 0.4875 0.4928 0.8518 0.6243 0.4964 0.7344 0.7020 0.8146 0.7541 0.8223
CodeT5-base 0.5000 0.5000 0.9143 0.6465 0.4974 0.6096 0.5714 0.8770 0.6920 0.7046
CodeT5p-220M 0.5036 0.5023 0.7661 0.6068 0.5109 0.6533 0.6276 0.7540 0.6850 0.7296
CodeT5p-770M 0.4955 0.4660 0.4298 0.4471 0.4747 0.5407 0.5102 0.7756 0.6159 0.6043
NatGen 0.4973 0.4951 0.2696 0.3491 0.5038 0.6312 0.6043 0.7607 0.6735 0.7213
Deepseek-Coder-1.3B-base 0.4719 0.4723 0.8385 0.6042 0.4867 0.5482 0.5250 0.7673 0.6235 0.5965
Granite-3B-Code-base-2k 0.4669 0.4656 0.6336 0.5368 0.4776 0.5075 0.4894 0.5584 0.5216 0.5587
StarCoder2-3B 0.4812 0.4814 0.8315 0.6098 0.4987 0.5696 0.5443 0.7197 0.6198 0.6049
CodeQWen1.5-7B 0.4901 0.4858 0.7838 0.5998 0.4919 0.5280 0.5048 0.9591 0.6615 0.5471
Qwen2.5_Coder-7B 0.4767 0.4705 0.5860 0.5220 0.4838 0.5468 0.5300 0.5064 0.5180 0.5915
CodeLlama-7b-hf 0.4994 0.5091 0.5489 0.5283 0.4843 0.5160 0.5242 0.5675 0.5450 0.5402
@@ 4.5.1. The general impact of code comments on fine-tuned CodeLMs performance
Compared to the fine-tuning results without code comments, most CodeLMs showed an overall performance improvement after fine-tuning with commented data, indicating that code comments provide additional contextual information, helping the model better identify potential CLBs. For UniXcoder-base, all metrics were improved after the inclusion of code comments, suggesting that the addition of code comments significantly enhanced the model’s performance in CLB detection, especially in terms of accuracy, F1 score, and AUC. GraphCodeBERT-base, after fine-tuning with code comments, saw an improvement in AUC from 0.7447 to 0.7898 and in precision from 0.6128 to 0.6304. Although recall dropped slightly (from 0.8241 to 0.7754), the increases in AUC and precision indicate that the inclusion of code comments helped improve the model’s performance of CLB detection. CodeT5-base, on the other hand, saw a substantial increase in recall from 0.6840 to 0.8770 and a moderate improvement in F1 score from 0.6586 to 0.6920. While accuracy and precision decreased, this indicates that code comments helped the model enhance its overall CLB detection capabilities. However, this improvement came at the cost of a slight reduction in precision. For CodeT5p-220M, fine-tuning with code comments resulted in a noticeable boost across several metrics: accuracy improved from 0.6364 to 0.6533, precision from 0.6148 to 0.6276, recall from 0.7307 to 0.7540, and F1 score from 0.6678 to 0.6850. This result shows that code comments contributed to an overall performance improvement for this model.
@@ 4.5.2. Impact of code comments on recall and precision
Code comments significantly improve recall, which is crucial for CLB detection tasks, as increasing recall helps reduce false negatives and capture more potential bugs. However, in some CodeLMs, this often comes with a decrease in precision. For example: The precision of CodeT5-base notably dropped from 0.6350 to 0.5714, despite a substantial increase in recall (from 0.6840 to 0.8770). This phenomenon suggests that the inclusion of code comments might lead to more false positives, as the boost in recall is accompanied by a significant decline in precision. Compared with the fine-tuning results without code comments, Granite-3B-Code-base-2k’s accuracy dropped from 0.5516 to 0.5075, and precision decreased from 0.5275 to 0.4894. Despite an increase in recall from 0.6951 to 0.7197, the decline in precision and F1 score (from 0.5998 to 0.5584) indicate that the inclusion of code comments did not result in an overall performance improvement. For Deepseek-Coder-1.3B-base, after fine-tuning with code comments, the recall increased from 0.7063 to 0.7673, and the F1 score rose from 0.6026 to 0.6235. Although both the recall and F1 score were improved, the AUC dropped from 0.6125 to 0.5965, with the accuracy and precision showing little change. This suggests that code comments may have led to more misclassifications, reducing the CodeLM’s discriminative ability.
Finding 4: The inclusion of code comments in the CLB dataset has indeed improved the performance of most CodeLMs under consideration, particularly in terms of recall and F1 score, which helps the CodeLMs capture more potential bugs in the CLB detection task. However, the inclusion of code comments may also lead to a decline in precision in CLB detection, especially when there is a trade-off between precision and recall. For some CodeLMs, fine-tuning on the CLB dataset with code comments resulted in a significant drop in precision, without a sufficient increase in recall to compensate for this decline, ultimately impairing the overall performance of these CodeLMs.

### v2

@@ 4.2.2. Performance of Small and Large CodeLMs
As shown in the results of Table 8, the impact of fine-tuning varies across different CodeLMs. We further explored the differences in fine-tuning effects based on the CodeLM’s size and complexity. According to the definition in (Zhang et al., 2026), models with fewer than 1B parameters are categorized as small LMs, while models with 1B parameters or more are considered as large LMs.
@@
Small CodeLMs: Small-scale CodeLMs, such as UniXcoder-base, GraphCodeBERT-base, CodeBERT-base, CodeT5p-220M, and NatGen, exhibited strong overall performance. In particular, UniXcoder-base demonstrated the best results across almost all metrics, suggesting that small CodeLMs can gain more knowledge transfer and performance enhancement from fine-tuning on the CLB dataset. On the other hand, among the small CodeLMs, CodeT5p-770M, with 770M parameters, showed the poorest overall performance. Despite having the best recall metric among all models (0.8779), it performed worse in the other four metrics. The result suggests that, for the CLB detection task, large CodeLMs do not necessarily outperform small CodeLMs in terms of overall performance.
Large CodeLMs: Among the large CodeLMs, those with 3B or fewer parameters, specifically Deepseek-Coder-1.3B-base, Granite-3B-Code-base-2k, and StarCoder2-3B, showed moderate performance, yet still outperformed the baseline model GPT-4o-mini on the CLB detection task. As for the remaining three 7B large models — CodeQWen1.5-7B, Qwen2.5_Coder-7B, and CodeLlama-7b-hf — their performance was relatively limited compared to the smaller models, only slightly surpassing GPT-4o-mini on certain metrics while falling behind it on others.
Finding 1: Among the 13 CodeLMs, UniXcoder-base and GraphCodeBERT-base demonstrate strong performance, with UniXcoder-base achieving the best overall results across all metrics. The performance of the 7B models, such as Qwen2.5_Coder-7B and CodeLlama-7b-hf are relatively limited. Overall, within the scope of our experimental setup, small CodeLMs tend to achieve better performance than large ones in the CLB detection task. The inference-only GPT-4o-mini baseline is consistently outperformed by fine-tuned CodeLMs, highlighting the importance of task-specific adaptation.
4.3. RQ2: Impact of Different Fine-Tuning Datasets on CLB Detection Performance
To answer RQ2, we further investigate whether CodeLMs fine-tuned on single-language bug datasets can effectively improve their ability to detect CLB. Specifically, five CodeLMs that perform well in RQ1 are selected. These CodeLMs are fine-tuned on 2 representative single-language bug datasets, namely CodeNet and CVEfixes, as well as on the CLB dataset constructed in this work. After fine-tuning, all CodeLMs are evaluated on the CLB test set to ensure a fair comparison of different fine-tuning strategies in the CLB detection task.
Tables 9 and 10 report the experimental results. When the CodeLMs are fine-tuned only on single-language bug datasets, either CodeNet or CVEfixes, their detection performance on the CLB dataset is generally limited. In particular, the accuracy and precision of most CodeLMs are close to 0.5, and the AUC values also remain around 0.5. This indicates that the CodeLMs achieve near random performance in the CLB detection task. Although some CodeLMs achieve relatively higher recall values and slightly improved F1 scores, this improvement is limited and is not reflected in the AUC metric. This result suggests that the CodeLMs fail to achieve an effective balance between precision and recall in CLB detection.
@@ 4.3. RQ2: Impact of Different Fine-Tuning Datasets on CLB Detection Performance
Table 9. CLB Detection Performance of CodeLMs Fine-Tuned on CodeNet and the CLB Dataset
CodeLM Fine-tuned in CodeNet Fine-tuned in CLB Dataset
@@
CodeBERT-base 0.5009 0.5006 0.7253 0.5924 0.4869 0.6331 0.6022 0.7842 0.6813 0.7131
GraphCodeBERT-base 0.5108 0.5081 0.6732 0.5792 0.5012 0.6517 0.6128 0.8241 0.7029 0.7447
UniXcoder-base 0.4973 0.4982 0.7343 0.5936 0.5028 0.7172 0.6839 0.8079 0.7407 0.8088
CodeT5p-220M 0.5162 0.5105 0.7864 0.6191 0.5072 0.6364 0.6148 0.7307 0.6678 0.7313
NatGen 0.4982 0.4987 0.6934 0.5802 0.5010 0.6364 0.6095 0.7594 0.6763 0.7278
Table 10. CLB Detection Performance of CodeLMs Fine-Tuned on CVEfixes and the CLB Dataset
CodeLM Fine-tuned in CVEfixes Fine-tuned in CLB Dataset
@@
CodeBERT-base 0.5036 0.5022 0.8366 0.6276 0.5024 0.6331 0.6022 0.7842 0.6813 0.7131
GraphCodeBERT-base 0.4937 0.4959 0.7684 0.6028 0.4999 0.6517 0.6128 0.8241 0.7029 0.7447
UniXcoder-base 0.5018 0.5011 0.7864 0.6122 0.4901 0.7172 0.6839 0.8079 0.7407 0.8088
CodeT5p-220M 0.5018 0.5012 0.7504 0.6010 0.5053 0.6364 0.6148 0.7307 0.6678 0.7313
NatGen 0.5152 0.5094 0.8217 0.6289 0.4938 0.6364 0.6095 0.7594 0.6763 0.7278
In contrast, when the CodeLMs are directly fine-tuned on the CLB dataset, their performance in CLB detection improves significantly. All CodeLMs show consistent improvements in accuracy, precision, F1 score, and AUC. For example, UniXcoder-base achieves an F1 score of 0.7407 and an AUC of 0.8088 after fine-tuning on the CLB dataset, which is clearly better than its performance after fine-tuning on single-language datasets. Similar performance gains are observed for the other CodeLMs, indicating that fine-tuning on CLB data provides a stable and effective improvement for CLB detection.
Overall, these results indicate that fine-tuning only on single-language bug datasets does not transfer well to the CLB detection task. CLB usually arise from interactions and dependencies between different PLs, and their characteristics are fundamentally different from those of single-language bugs. Therefore, CodeLMs can effectively learn relevant patterns and achieve reliable detection when they are fine-tuned on CLB datasets that explicitly contain cross-language semantics and interaction information.
Finding 2: CodeLMs fine-tuned on single-language bug datasets such as CodeNet and CVEfixes show limited performance in the CLB detection task, with overall results close to random performance. In contrast, fine-tuning on the CLB dataset significantly improves detection performance. This finding indicates that CLB detection requires dedicated datasets and training strategies, and single-language bug data alone cannot adequately support this task.
@@ 4.4.2. Impact of the Token Sequence Length
Finding 3: Increasing the size of the fine-tuning dataset for CodeLMs can indeed improve the model’s performance in CLB detection to some extent. However, within the scope of our experimental setup, increasing the token sequence length does not necessarily lead to better performance. Different CodeLMs show varying sensitivity to token sequence length; some CodeLMs (e.g., CodeT5p-220M) perform exceptionally well with shorter token sequence lengths, and their performance decreases as the token sequence length increases. On the other hand, some CodeLMs (e.g., GraphCodeBERT-base) perform better with longer token sequence lengths.
@@ 4.5. RQ4: Impact of Code Comments on Fine-tuned CodeLM Performance
We have already conducted fine-tuning experiments and analysis on the CLB dataset without code comments, and the results are shown in Table 8. To answer RQ4, we carried out fine-tuning experiments on the CLB dataset with code comments, and the results are shown in Table 11. In addition, we include GPT-4o-mini as an inference-only reference baseline evaluated on the same commented test set using prompt-based reasoning, without any task-specific fine-tuning or parameter updates. By comparing the performance of CodeLMs fine-tuned on the CLB dataset with and without comments, we analyze the impact of code comments on the performance of CodeLMs in the following two aspects.
Table 11. Performance of CodeLMs in detecting CLBs after fine-tuning on the CLB dataset with code comments
CodeLM Acc Prec Rec F1 AUC
GPT-4o-mini (Baseline) 0.5152 0.5171 0.4589 0.4863 0.5152
CodeBERT-base 0.6304 0.6008 0.7768 0.6776 0.7108
GraphCodeBERT-base 0.6604 0.6304 0.7754 0.6954 0.7898
UniXcoder-base 0.7344 0.7020 0.8146 0.7541 0.8223
CodeT5-base 0.6096 0.5714 0.8770 0.6920 0.7046
CodeT5p-220M 0.6533 0.6276 0.7540 0.6850 0.7296
CodeT5p-770M 0.5407 0.5102 0.7756 0.6159 0.6043
NatGen 0.6312 0.6043 0.7607 0.6735 0.7213
Deepseek-Coder-1.3B-base 0.5482 0.5250 0.7673 0.6235 0.5965
Granite-3B-Code-base-2k 0.5075 0.4894 0.5584 0.5216 0.5587
StarCoder2-3B 0.5696 0.5443 0.7197 0.6198 0.6049
CodeQWen1.5-7B 0.5280 0.5048 0.9591 0.6615 0.5471
Qwen2.5_Coder-7B 0.5468 0.5300 0.5064 0.5180 0.5915
CodeLlama-7b-hf 0.5160 0.5242 0.5675 0.5450 0.5402
@@ 4.5.1. The general impact of code comments on fine-tuned CodeLMs performance
Compared to the fine-tuning results without code comments, most CodeLMs showed an overall performance improvement after fine-tuning with commented data, indicating that code comments provide additional contextual information, helping the model better identify potential CLBs. For UniXcoder-base, all metrics were improved after the inclusion of code comments, suggesting that the addition of code comments significantly enhanced the model’s performance in CLB detection, especially in terms of accuracy, F1 score, and AUC. GraphCodeBERT-base, after fine-tuning with code comments, saw an improvement in AUC from 0.7447 to 0.7898 and in precision from 0.6128 to 0.6304. Although recall dropped slightly (from 0.8241 to 0.7754), the increases in AUC and precision indicate that the inclusion of code comments helped improve the model’s performance of CLB detection. CodeT5-base, on the other hand, saw a substantial increase in recall from 0.6840 to 0.8770 and a moderate improvement in F1 score from 0.6586 to 0.6920. While accuracy and precision decreased, this indicates that code comments helped the model enhance its overall CLB detection capabilities. However, this improvement came at the cost of a slight reduction in precision. For CodeT5p-220M, fine-tuning with code comments resulted in a noticeable boost across several metrics: accuracy improved from 0.6364 to 0.6533, precision from 0.6148 to 0.6276, recall from 0.7307 to 0.7540, and F1 score from 0.6678 to 0.6850. This result shows that code comments contributed to an overall performance improvement for this model. For reference, GPT-4o-mini achieves moderate performance on the commented CLB dataset under inference-only settings. However, its overall performance remains consistently lower than that of fine-tuned CodeLMs, indicating that the availability of code comments alone is insufficient to fully compensate for the absence of task-specific adaptation.
@@ 4.5.2. Impact of code comments on recall and precision
Code comments significantly improve recall, which is crucial for CLB detection tasks, as increasing recall helps reduce false negatives and capture more potential bugs. However, in some CodeLMs, this often comes with a decrease in precision. For example: The precision of CodeT5-base notably dropped from 0.6350 to 0.5714, despite a substantial increase in recall (from 0.6840 to 0.8770). This phenomenon suggests that the inclusion of code comments might lead to more false positives, as the boost in recall is accompanied by a significant decline in precision. Compared with the fine-tuning results without code comments, Granite-3B-Code-base-2k’s accuracy dropped from 0.5516 to 0.5075, and precision decreased from 0.5275 to 0.4894. Despite an increase in recall from 0.6951 to 0.7197, the decline in precision and F1 score (from 0.5998 to 0.5584) indicate that the inclusion of code comments did not result in an overall performance improvement. For Deepseek-Coder-1.3B-base, after fine-tuning with code comments, the recall increased from 0.7063 to 0.7673, and the F1 score rose from 0.6026 to 0.6235. Although both the recall and F1 score were improved, the AUC dropped from 0.6125 to 0.5965, with the accuracy and precision showing little change. This suggests that code comments may have led to more misclassifications, reducing the CodeLM’s discriminative ability. A similar tendency can be observed from the inference-only GPT-4o-mini baseline, whose recall does not substantially benefit from the presence of code comments. This observation further suggests that effectively leveraging code comments for CLB detection requires model adaptation through fine-tuning, rather than relying solely on prompt-based reasoning.
Finding 4: Incorporating code comments into the CLB dataset generally improves CodeLM performance, particularly in recall and F1 score, enabling the detection of more potential CLBs. However, such improvements may come at the cost of reduced precision for some models, leading to limited or even negative overall performance gains. Moreover, results from the inference-only GPT-4o-mini baseline suggest that code comments alone are insufficient to guarantee improved CLB detection performance, underscoring the necessity of task-specific fine-tuning to effectively leverage comment-level information.